THE PERMANENT ELECTRIC DIPOLE MOMENTS OF THE ^{3}Π and ^{3}Δ STATES OF RUTHENIUM MONOCARBIDE, RuC

\underline{WILTON L. VIRGO}, TIMOTHY C. STEIMLE AND LAURA E. AUCOIN, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287-1604; JOHN M. BROWN, Physical and Theoretical Chemistry Laboratory, Oxford University, South Parks Road, Oxford, England OX1 3QZ.

Elucidating the structure and function of the chemically pervasive transition metal-carbon bond is a problem of both fundamental and applied scientific interest. Recent insight into the nature of the metal-carbon bond has been due to high-resolution gas-phase spectroscopy of the diatomic monocarbides, where measurement of permanent electric dipole moments and hyperfine structure are particularly informative. Ruthenium monocarbide has garnered recent interest^{a} ^{b} ^{c} d ^{e} ^{f} due to its ease of production, intense visible electronic transitions and large magnetic hyperfine structure. We report on the investigation of the (0,0) {12.7}^{3}Π_{2} −{0.1}^{3}Δ_{3} and (0,0) {13.9}^{3}Π_{1} −{0.9}^{3}Δ_{2} band systems using high-resolution laser induced fluorescence spectroscopy. Stark shifts of the ^{102}RuC and ^{104}RuC isotopomers were analyzed to produce the magnitude of the permanent electric dipole moments for the ^{3}Δ_{3},^{3} Δ_{2},^{3} Π_{2} and ^{3}Π_{1} states. The measured moments prompt a discussion of the electronic structure and bonding in the ^{3}Δ and ^{3}Π electronic states. Dipole moment trends are most informative when coupled to molecular orbital correlation diagrams. The measured dipole moments can be used as true benchmarks for rigorous electronic structure calculations. A comparison with isovalent FeC is made. ^{g}

aT. C. Steimle and W. L. Virgo J. Chem. Phys. \underline{118}, 2620 (2003).

bI. Shim and K. A. Gingerich Chem. Phys. Lett. \underline{317}, 338 (2000).

cJ. D. Langenberg, R. S. DaBell, L. Shao, D. Dreessen and M. Morse J. Chem. Phys. \underline{109}, 7863 (1998).

dR. S. DaBell, R. G. Meyer and M. D. Morse J. Chem. Phys. \underline{114}, 2938 (1998).

eI. Shim, H. C. Finkbeiner and K. A. Gingerich J. Phys. Chem. \underline{91}, 3171 (1987).

fR. Scullman and B. Thelin Phys. Scr. \underline{3}, 19 (1971).

gT. C. Steimle, W. L. Virgo and D. A. Hostutler J. Chem. Phys. \underline{117}, 1511 (2002).